# Disentanglement in Implicit Causal Models via Switch Variable

### v1

@@ 3.1 Structural Causal Models for Intervention Data
(Structural Causal Models) A structural causal model (SCM) is a tuple $\mathcal{C}=(\mathcal{F},\mathcal{Z},\mathcal{E},\mathcal{G})$ with the following components:
  • The domain of causal variables $\mathcal{Z}=\mathcal{Z}_{1}\times\mathcal{Z}_{2}\times\ldots\times\mathcal{Z}_{n}$.
  • The domain of exogenous variables $\mathcal{E}=\mathcal{E}_{1}\times\mathcal{E}_{2}\times\ldots\times\mathcal{E}_{n}$.
  • a directed acyclic graph $\mathcal{G}(\mathcal{C})$ over the causal and exogenous variables.
  • causal mechanism $f_{i}\in\mathcal{F}$ which maps an assignment of parent values for the parents $\mathcal{Z}_{pa_{i}}$ plus an exogenous variable value for $\mathcal{E}_{i}$ to a value of causal variable $Z_{i}$.
A decoder function $g(\mathbf{z})=\mathbf{x}$ maps a vector of causal values to observed values. The causal variables $\mathcal{Z}$ are unobserved and the goal is to infer them from interventional data. For each causal variable, a solution function $s_{i}:\mathcal{E}\rightarrow\mathcal{Z}_{i}$ deterministically maps the exogenous variables to a value for $\mathcal{Z}_{i}$. In implicit modeling, we learn the solution functions $s_{i}$ directly, rather than defining them through local mechanisms $f_{i}$. We write $s$ for the set of all solution functions $s_{i}$, so $s(\mathbf{e})=\mathbf{z}.$ Since exogenous variables are assumed to be mutually independent in a causal model, we can conveniently factorize $p(\mathcal{E})=\Pi_{i}p(\mathcal{E}_{i})$ whereas the priors of causal variables should be factorized according to the Causal Markov condition Schölkopf et al. (2021) i.e., $p(\mathcal{Z})=\Pi_{i}p(\mathcal{Z}_{i}|\mathcal{Z}_{pa_{i}})$.
Identifying causal models from data can be complex and is often studied within classes of models such as those identifiable up to affine transformations. For example, in the context of nonlinear Independent Component Analysis (ICA), the generative process also involves a mixture function $g$ of latent causal variables $\mathcal{Z}\in\mathbb{R}^{n}$, resulting in observations $\mathcal{X}\in\mathbb{R}^{n}$ Lachapelle et al. (2022); Zheng et al. (2022). However, a significant distinction between causal representation learning and nonlinear-ICA is that in the former, the causal variables $\mathcal{Z}$ may have complex dependencies. Our objective in this paper is to recover $\mathcal{E}$ from $\mathcal{X}$ and eventually map $\mathcal{E}$ to $\mathcal{Z}$ using solution functions.
Identifying a causal model from observational data is not trivial and requires assumptions on the parameters of the model Ahuja et al. (2023). Adding information about interventions in addition to observations, helps to identify causal variables by exhibiting the effect of changing a causal variable on the observed variables. An interventional data point $(x,\tilde{x},i)$ includes the pre-intervention observation $x$, the post-intervention observation $\tilde{x}$, and intervention target $i\in\mathcal{I}$ where $\mathcal{I}$ is the set of intervention targets. The post-intervention data is generated by a soft intervention that targets one of the causal variables $\mathcal{Z}$. A soft intervention can be represented with causal mechanism switch variable $\mathcal{V}\in\mathbb{R}^{n}$. Adding switch variables leads to the concept of an implicit augmented SCM.
@@
(Implicit Augmented Structural Causal Models) The Augmented Structural causal models (ASCMs) are defined as $\mathcal{A}=(\mathcal{F},\mathcal{Z},\mathcal{E},\mathcal{G},\mathcal{V})$ where $\mathcal{V}\in\mathbb{R}^{n}$ is the causal mechanism switch variable which models the effect of soft interventions on causal mechanisms $\mathcal{F}$:
$\forall i,\>\tilde{\mathcal{Z}_{i}}=\tilde{s_{i}}(\tilde{\mathcal{E}}_{/i},\tilde{\mathcal{E}}_{i})=s_{i}(\mathcal{E}_{/i},\mathcal{E}_{i},\mathcal{V}),$
where $\tilde{s_{i}}$ is the new solution function resulted from the soft intervention, $\tilde{\mathcal{Z}}_{pa_{i}}$ is the altered set of parents due to intervention, and ~$\tilde{\mathcal{E}}_{i}$ is the post-intervention exogenous variable.
Causal mechanism switch variable:
The intuition behind using $\mathcal{V}$ is to separate the effect of soft intervention into two: (1) The effect on causal mechanisms and parents (2) The effect on exogenous variable. Let’s consider a set of images containing objects where an action is performed on them (see images in Figure A5). We can say that causal variables in these images are the objects’ attributes such as shape, color, and size and the actions change these attributes. Furthermore, it can be asserted that the camera angle within a given image may influence the shape of the object. This could be conceptualized as a shift in the conditional distribution of object attributes conditioned on their respective parent elements, resulting from various actions. Consequently, the images can be seen as a result of soft interventions where an action changes the intrinsic attributes of an object but also changes the causal variable’s conditional distribution. In this case, if we had a knowledge of how the camera angle is changed, then we could separate the effect of soft intervention. In other words, if $\mathcal{V}$ is observed, then we can extract the effect of the intervention that we are interested in (i.e., the effect on the causal variable itself). Observability of $\mathcal{V}$ is given as an assumption below:
The new parent set of $\mathcal{Z}_{i}$, post intervention, will be $\widetilde{PA}_{i}=PA_{i}\cup\mathcal{V}$ which is related to $\mathcal{Z}_{i}$ by the following conditional probability:
$p(z_{i}|\widetilde{pa}_{i})=\begin{cases}p(z_{i}|pa_{i},v)\quad&\text{if $z_{i}$ is intervened softly.}\\ p(z_{i}|pa_{i})\quad&\text{if no interventions}.\end{cases}$ (5)
In Appendix A2, we discuss how $\mathcal{V}$ can also handle the change in the set of parents due to a soft intervention. The usage of $\mathcal{V}$ in soft interventions is analogous to augmented networks in Pearl (2005) which were mainly designed for hard interventions. Pearl even foresaw this possibility in the original paper, saying "One advantage of the augmented network representation is that it is applicable to any change in the functional relationship $f_{i}$ and not merely to the replacement of $f_{i}$ by a constant."Pearl (2005)
3.2 Data Generating Process
The data generation process for an implicit SCM can be defined as follows.
Definition 3.3 (Data Generation Process).
Consider a latent causal model where the underlying SCM is defined by a continuous latent space $\mathcal{E}\in\mathbb{R}^{n}$ with independent probabilities $p(\mathcal{E}_{i})$ which admits a solution function $s$. The pre-intervention data is generated as:
$\begin{split}&e\sim p(\mathcal{E}),\quad z=s(e),\quad x=g(z).\\ &\tilde{e}\sim p(\tilde{\mathcal{E}}),\quad\tilde{z}=\tilde{s}(\tilde{e}),\quad\tilde{x}=g(\tilde{z}).\end{split}$ (6)
Figure 1:  General overview of the SoftCD for learning an encoder, decoder, and solution function from interventional data. FC represents a fully connected neural network, and S the solution function.
Assumption 3.4.
(Identical correspondence assumptions)
  1. 1.
    Atomic Interventions: For every paired sample $(x,\tilde{x})$, only one causal variable is targeted by an intervention.
  2. 2.
    Known Targets: Targets of soft interventions are known.
  3. 3.
    Post-intervention Exogenous Variables: The exogenous value changes only for the intervened causal variable, while the others maintain their pre-intervention values.
@@ 3.2 Data Generating Process
  4. 4.
    Sufficient Variability: Soft interventions alter causal mechanisms to introduce sufficient variability Lachapelle et al. (2022). These interventions should modify causal mechanisms to ensure non-overlapping conditional distributions of causal variables (refer to Figure A2).
  5. 5.
    Diffeomorphic decoder and causal mechanisms: Diffeomorphism guarantees no information loss and avoids abrupt changes in the function’s image.
Without these assumptions, observed changes in $\tilde{x}$ become ambiguous, as multiple interventions or alterations in exogenous variables may result in overlapping effects on the observed variables. For instance, when examining the combined impact of both heater and lights on room temperature, changing them simultaneously would obscure the precise influence of each on the room temperature.
3.3 Theoretical Identifiability Analysis for Implicit SCMs with Soft Interventions
Since our focus in this paper is to disentangle the causal variables using soft interventions, first we define causal disentanglement and identifiability up to reparameterization.
Definition 3.5.
(Component-wise Transformation) Let $\phi$ be a transformation (1-1 onto mapping) between product spaces $\phi:\Pi_{i=1}^{n}\mathcal{X}_{i}\rightarrow\Pi_{i=1}^{n}\mathcal{Y}_{i}$. If there exist local transformations $\phi_{i}$ such that $\forall i,\>\forall x,\>\phi(x_{1},x_{2},...,x_{n})_{i}=\phi_{i}(x_{i})$, then $\phi$ is a component-wise transformation.
Definition 3.6.
(Equivalence up to component-wise reparameterization) Let $\mathcal{M}=(\mathcal{A},\mathcal{X},g,\mathcal{I})$ and $\mathcal{M}^{\prime}=(\mathcal{A}^{\prime},\mathcal{X},g^{\prime},\mathcal{I})$ be two Latent Causal Models (LCM) based on ASCMs $\mathcal{A},\mathcal{A}^{\prime}$ with shared observation space $\mathcal{X}$, shared causal variables $\mathcal{Z}$, shared intervention targets $\mathcal{I}$, and respective decoders $g$ and $g^{\prime}$.
We say that $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are equivalent up to component-wise reparameterization $\mathcal{M}\sim_{r}\mathcal{M}^{\prime}$ if there exists a component-wise transformation $\phi_{\mathcal{Z}}$ from the causal variables $\mathcal{Z}$ to the causal variables $\mathcal{Z}^{\prime}$ and a component-wise transformation $\phi_{\mathcal{E}}$ between $\mathcal{E}$ and $\mathcal{E}^{\prime}$ such that:
  1. 1.
    Indices are preserved (i.e., $\phi_{i}(z_{i})=z^{\prime}_{i}$ and $\phi_{i}(e_{i})=e^{\prime}_{i}$). Corresponding edges are preserved (i.e., $Z_{i}\rightarrow Z_{j}$ holds in $G$ iff $Z^{\prime}_{i}\rightarrow Z^{\prime}_{j}$ holds in $G^{\prime}$. Edges $E_{i}\rightarrow Z_{i}$ should be preserved as well.)
  2. 2.
    The exogenous transformation preserves the probability measure on exogenous variables $p_{\mathcal{E}^{\prime}}=(\phi_{\mathcal{E}})_{*}p_{\mathcal{E}}$ (Definition A2.2).
  3. 3.
    The causal transformation preserves the probability measure on causal variables $p_{\mathcal{Z}^{\prime}}=(\phi_{\mathcal{Z}})_{*}p_{\mathcal{Z}}$ (Definition A2.2).
Definition 3.7 (Disentanglement Lachapelle et al. (2022)).
Given a groundtruth LCM $\mathcal{M}^{*}$, we say a learned LCM $\mathcal{M}$ is disentangled when $\mathcal{M}^{*}$ and $\mathcal{M}$ are equivalent up to reparameterization.
Figure 2: Causal graph models in the presence of Hard (a) and Soft (b) interventions.
Now we can analyze how and where soft interventions violate identifiability in models that disentangle causal variables using hard interventions.
Soft intervention:
Let’s consider two SCMs $\mathcal{C}$ and $\mathcal{C}^{\prime}$ and use $o$ to denote all variables except $i$. The string diagrams for soft and hard interventions in $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are depicted in Figure 2. The major difference of soft intervention (Figure 2(b)) with hard intervention (Figure 2(a)) is that $\tilde{\mathcal{Z}_{i}}$ is no longer disconnected from its parents and its causal mechanism $f_{i}$ is affected by the intervention. Thus, with a hard intervention, we know the post-intervention parents of a node $Z_{i}$ (there are none), whereas with soft interventions, the parents themselves may not change. This is why identifying the causal mechanisms is more difficult for soft interventions. Soft intervention data yield fewer constraints on the causal graph structure than hard intervention data.
In order for the causal model to be identifiable according to Definition 3.6, $\mathcal{Z}^{\prime}_{i}$ should only be the function of its corresponding $\mathcal{Z}_{i}$ and not $\mathcal{Z}_{o}$. Mathematically, we say that $\phi$ should be a component-wise transformation according to Definition 3.5. This characteristic will be violated in the soft intervention case and $\tilde{\mathcal{Z}^{\prime}_{i^{\prime}}}$ will no longer be a function of $\tilde{\mathcal{Z}_{i}}$ alone. To address this and model the effect of soft interventions, we introduce the causal mechanism switch variable $\mathcal{V}$. As soft interventions can also alter the set of parents of a causal variable, we propose to use a modulated form of $\mathcal{V}$ to model the soft intervention effects on each variable as an additive effect with a nonlinear function $h_{i}$ such that:
$\forall i,\>\tilde{\mathcal{Z}_{i}}=\tilde{s_{i}}(\tilde{\mathcal{E}}_{/i},\tilde{\mathcal{E}}_{i})=s_{i}(\mathcal{E}_{/i},\tilde{\mathcal{E}}_{i},\mathcal{V})\approx s_{i}(\mathcal{E}_{/i},\tilde{\mathcal{E}}_{i})+h_{i}(\mathcal{V}).$ (7)
As the parental set for each causal variable is not known, the inclusion of $\mathcal{V}$ in Equation 7 enables the model to encompass variations in the parental sets of all causal variables in $\mathcal{V}$. Therefore, there is a switch variable $\mathcal{V}_{i}$ for each causal variable $\mathcal{Z}_{i}$. Subsequently, it can modulate the impact of a soft intervention on the intervened causal variable using $h_{i}$.
Assumption 3.8.
(Observability of modulated $\mathcal{V}$) Given an intervention sample $(x,\tilde{x},i)$ and linear decoders, we can approximate the soft intervention effects $h_{i}(\mathcal{V})$ as follows:
[…]
$\tilde{x}-x=g(\tilde{z})-g(z)\approx g(\tilde{z}-z)=g(\Delta{e_{A}}-h(v)),$
where $h(v)$ and $\Delta{e_{A}}$ are the vectors indicating the change in causal mechanism and change in effect of ancestral exogenous variables of each causal variable, respectively. Note that elements of $h(v)$ will be all zero except for the intervened causal variable. Consequently, with linear mixing functions and some pre-processing on observed samples (here subtraction), we can observe $h_{i}(v)$.
Theorem 3.9.
(Identifiability of augmented latent causal models. The proof is given in Appendix A2.) Let $\mathcal{M}=(\mathcal{A},\mathcal{X},g,\mathcal{I})$ and $\mathcal{M}^{\prime}=(\mathcal{A}^{\prime},\mathcal{X},g^{\prime},\mathcal{I})$ be two LCMs with shared observation space $\mathcal{X}$ and shared intervention targets $\mathcal{I}$. Suppose the following conditions are satisfied:
  1. 1.
    Identical correspondence assumptions explained in 3.4.
  2. 2.
    Soft interventions satisfy Assumption 3.8.
  3. 3.
    The causal and exogenous variables are real-valued.
Then the following statements are equivalent:
The proof is in Appendix A2.
3.4 Training Objective
Consequently, there will be three latent variables in SoftCD:
  1. 1.
    A causal mechanism switch variable $\mathcal{V}$.
  2. 2.
    The pre-intervention causal variables $\mathcal{E}$.
  3. 3.
    The post-intervention causal variables $\tilde{\mathcal{E}}$.
$\log p(x,\tilde{x})\geq$ $E_{q(e,\tilde{e},v|x,\tilde{x})}\Bigl{[}\log p(x,\tilde{x}|e,\tilde{e},v)\Big{]}-KLD(q(e,\tilde{e},v|x,\tilde{x})||p(e,\tilde{e},x))$ (8)
$=E_{q(v|\tilde{x}-x)\cdot q(e|x)\cdot q(\tilde{e}|\tilde{x})}\Bigl{[}\log(p(x|e)p(\tilde{x}|\tilde{e})p(\tilde{x}-x|v))\Big{]}$
$-KLD(q(v|\tilde{x}-x)\cdot q(e|x)\cdot q(\tilde{e}|\tilde{x})||p(\tilde{e}|e,v)p(v)p(e))$
@@ 3.4 Training Objective
  1. 1.
    $p(e)$ is the prior distribution over exogenous variables $e$.
  2. 2.
    $p(v)$ is the prior distribution over switch variables $v$.
  3. 3.
    $p(\tilde{e}|e,v)$ is a transition model that shows how the exogeneous variables change as a function of the intervention.
We factorize the posterior with a mean-field approximation $q(v,e,\tilde{e}|x,\tilde{x})=q(v|\tilde{x}-x)\cdot q(e|x)\cdot q(\tilde{e}|\tilde{x})$ and, following our data generation model (Figure A1), the reconstruction probability as $p(x,\tilde{x}|e,\tilde{e},v)=p(x|e)p(\tilde{x}|\tilde{e})p(\tilde{x}-x|v)$. The prior over latent variables is factorized as $p(\tilde{e},e,v)=p(\tilde{e}|e,v)p(v)p(e)$(Figure A1). As exogenous variables are mutually independent $p(e)=\Pi_{i}p(e_{i})$ and we assume $p(e_{i})$ and $p(v)$ to be standard Gaussian (See Appendix A2). Furthermore, as we assume $e_{i}=\tilde{e_{i}}$ for all non-intervened variables, the $p(\tilde{e}|e,v)$ will be as follows:
$p(\tilde{e}|e,v)=\Pi_{i\notin I}\delta(\tilde{e_{i}}-e_{i})\Pi_{i\in I}p(\tilde{e_{i}}|e_{i},v)=\Pi_{i\notin I}\delta(\tilde{e_{i}}-e_{i})\Pi_{i\in I}p(\tilde{z_{i}}|e_{i},v)\left|\frac{\partial\tilde{z_{i}}}{\partial\tilde{e_{i}}}\right|$ (9)
The last equality is obtained from the Change of Variable Rule in probability theory, applied to the solution function $\tilde{z}_{i}=\tilde{s_{i}}(\tilde{e}_{i};e,v)$ that maps the exogeneous variable $\tilde{e}_{i}$ to the causal variable $\tilde{z}_{i}$. We implement the solution function using a location-scale noise models Immer et al. (2023) as also practiced in Brehmer et al. (2022), which defines an invertible diffeomorphism. For simplicity, in our experiments, we are only going to change the $loc$ network in post-intervention. Therefore, $h_{i}(v)$ will be used as:
$\tilde{z}_{i}=\tilde{s_{i}}(\tilde{e}_{i};e,v)=\frac{\tilde{e}_{i}-(loc(e_{/i})+h_{i}(v))}{scale(e_{/i})},$ (10)
where $loc:\mathbb{R}^{n-1}\rightarrow\mathbb{R}$ and $scale:\mathbb{R}^{n-1}\rightarrow\mathbb{R}$ are fully connected networks calculating the first and second moments, respectively. We assume $p(\tilde{z}_{i}|e_{i},v)$ to be a Gaussian whose mean is determined by $e_{i}$ and $v$ as $e_{i}$ is a parent of $\tilde{z}_{i}$ and we do not have knowledge about the other ancestors in implicit modeling. We also obtain $v$ from $\tilde{x}-x$ to ensure that $v$ should capture the effects of soft intervention, since we assume the changes in $\tilde{x}$ are due to the soft intervention only. The general overview of the model is illustrated in Figure 1.

### v2

@@ 3.2 Causal Mechanisms Switch Variable
(Augmented Implicit Causal Models) An Augmented Implicit Causal Models (AICMs) is defined as $\mathcal{A}=(\mathcal{S},\mathcal{Z},\mathcal{E},\mathcal{V})$ where $\mathcal{V}\in\mathbb{R}^{n}$ is the causal mechanism switch variable which models the effect of soft interventions on solution functions $\mathcal{S}$:
$\forall i,\>\tilde{\mathcal{Z}_{i}}=\tilde{s}_{i}(\tilde{\mathcal{E}}_{i};\tilde{\mathcal{E}}_{/i})=s_{i}(\tilde{\mathcal{E}}_{i};\mathcal{E}_{/i},h_{i}(\mathcal{V})),$ (1)
where $\tilde{s_{i}}$ is the new solution function resulting from the soft intervention, ~$\tilde{\mathcal{E}}_{/i}$ is the altered set of all exogenous variables except $i$, including the ancestral exogenous variables, due to intervention, and ~$\tilde{\mathcal{E}}_{i}$ is the post-intervention exogenous variable.
The usage of $\mathcal{V}$ in soft interventions is analogous to augmented networks in [23] which were mainly designed for hard interventions. Pearl [23] even foresaw this possibility by saying:  "One advantage of the augmented network representation is that it is applicable to any change in the functional relationship $f_{i}$ and not merely to the replacement of $f_{i}$ by a constant."
By using Taylor’s expansion, we can expand the solution functions as follows:
[…]
where we’ll use $R_{i}$ as a short-hand for Equation 2. We define the separable dependence property for solution functions as $\exists h_{i}(v_{0}):s_{i}(\tilde{\mathcal{E}}_{i};\mathcal{E}_{/i},h_{i}(v_{0}))=s_{i}(\tilde{\mathcal{E}}_{i};\mathcal{E}_{/i})$. An example of such a scenario could be in location-scale noise models such as, $s_{i}(\tilde{e_{i}};e_{/i},h_{i}(v))=\tilde{e_{i}}+loc(e_{/i})+h_{i}(v)=\tilde{e_{i}}+loc(e_{/i})+v^{2}+v$ where $v_{0}$ would be zero . By assuming the separable dependence property, we can write the solution function in Equation 2 as:
$\begin{split}s_{i}(\tilde{\mathcal{E}}_{i};\mathcal{E}_{/i},h_{i}(\mathcal{V}))=s_{i}(\tilde{\mathcal{E}}_{i};\mathcal{E}_{/i})+R_{i}=s_{i}(\tilde{\mathcal{E}}_{i};\mathcal{E}_{/i})+\textit{soft intervention effect}\end{split}$ (3)
As a result, we can switch to pre-intervention solution functions. Subsequently, by modeling soft intervention effects using $h_{i}(\mathcal{V})$, we can recover pre-intervention solution functions. During inference, we simply disregard the $h_{i}(\mathcal{V})$ term in the solution functions. Nonetheless, it is possible to train the prior $p(\mathcal{V})$ to ensure that the separable dependence property is maintained for pre-intervention data.
Observability of switch variable
The intuition behind using $\mathcal{V}$ is to separate the effect of soft intervention on $\tilde{\mathcal{Z}}_{i}$ into two: (1) The effect on causal mechanisms and parents, and (2) The effect on exogenous variable $\mathcal{E}_{i}$. For example, we can say that causal variables in images of objects are the objects’ attributes such as shape, color, and size, and performing actions like "Fold" change these attributes. Furthermore, it can be asserted that the camera angle within a given image may influence the shape of the object. If the images were generated from a hard intervention, the camera angle remains fixed between pre and post intervention. However, the camera angle changes along with the performed actions indicating that the interventions are soft. In this case, if we had a knowledge of how the camera angle affects the attributes of objects, then we could separate the effect of soft intervention. In other words, if $\mathcal{V}$ is observed, then we can extract the effect of the intervention that we are interested in (i.e., the effect on the causal variable itself). For more details, refer to Figure A4.
Lacking an understanding of how soft intervention influences the causal model, a more complex model becomes necessary. Consequently, the term $R_{i}$ in Equation 2 would involve a higher order of $h_{i}(\mathcal{V})$. Therefore, we assume the observability of $\mathcal{V}$:
Assumption 3.3.
(Observability of $\mathcal{V}$) Given an intervention sample $(x,\tilde{x},i)$ and linear decoders, we can approximate the soft intervention effects $h_{i}(\mathcal{V})$ as follows:
[…]
where $R=[R_{0},R_{1},...,R_{n}]$ and $n$ is the number of causal variables. $R$ and $\Delta{e_{i}}$ are the vectors indicating the soft intervention effects and change in effect of the exogenous variable of the intervened causal variable, respectively. Note that elements of $R$ will be all zero except for the intervened causal variable. Consequently, with linear mixing functions and some pre-processing on observed samples (here subtraction), we can observe $R_{i}$.
Our synthetic data is generated using a linear decoder, however, the decoder for the real-world datasets is not necessarily linear. Therefore, we do not observe $\mathcal{V}$ from $\tilde{x}-x$ in the real-world dataset. Nevertheless, our findings suggest that incorporating soft interventions through $\mathcal{V}$ leads to superior performance compared to other implicit modeling approaches. Clearly, understanding the impact of soft interventions on the generative system of the dataset would result in improved outcomes.
(a) General overview of ICRL-SM
(b) Generative model
3.3 Identifiability Theorem for Implicit SCMs with Soft Interventions
In this paper, our focus lies in identifying the causal variables up to reparameterization through soft interventions. We first define identifiability up to reparameterization (Definition 3.4) and subsequently introduce the identifiability theorem 3.5. The proof of theorem is extensive and is available in full in Appendix A1.
We establish identifiability up to reparameterization, allowing for the mapping of causal variables $\mathcal{Z}$ and $\mathcal{Z}^{\prime}$ between two Latent Causal Models ($\mathcal{M}$ and $\mathcal{M}^{\prime}$) through component-wise transformations (Definition A1.2). Given our implicit modeling approach, lacking knowledge of the causal graph, we include all exogenous variables in the solution functions, as depicted in Equation 1. Notably, the causal graph remains unaltered during learning. To illustrate, we contrast hard interventions, which neglect parent influences, with soft interventions that acknowledge parental effects in a simple example. Consider a basic causal model $Z_{1}\rightarrow Z_{2}$ alongside a location-scale noise model [12] for the solution function, given by $\tilde{z}_{2}=\frac{\tilde{e}_{2}-\widetilde{loc}(e_{1})}{\widetilde{scale}(e_{1})}$. The distribution $p(\tilde{\mathcal{Z}}_{2})$ mean is $\frac{1}{\widetilde{scale}(e_{1})}\times\text{mean}(\tilde{\mathcal{E}}_{2})-\frac{\widetilde{loc}(e_{1})}{\widetilde{scale}(e_{1})}$ In the context of hard interventions, we can assume $p(\tilde{\mathcal{Z}}_{2}|\mathcal{Z}_{1})=p(\tilde{\mathcal{Z}}_{2})=N(0,1)$ as there are no parental effects. Consequently, the location and scale networks within the solution function tend to dampen parental effects, given the absence of parental influence in the ground-truth data. Contrarily, soft interventions exhibit parental influence in the ground-truth data, thus $p(\tilde{\mathcal{Z}}_{2}|\mathcal{Z}_{1})\neq N(0,1)$. Due to the lack of parental knowledge in implicit modeling, we model $p(\tilde{\mathcal{Z}}_{2}|\mathcal{Z}_{1})=p(\tilde{\mathcal{Z}}_{2}|\mathcal{E}_{2})$, as $\mathcal{E}_{2}$ is a known parent of $\tilde{\mathcal{Z}}_{2}$. Consequently, parental effects are propagated to $\mathcal{E}_{i}$ (the corresponding exogenous variable of each causal variable), violating identifiability up to reparameterization. By leveraging $\mathcal{V}$, we allow parental effects to propagate to $\mathcal{V}$ instead of $\mathcal{E}_{i}$.
@@ 3.3 Identifiability Theorem for Implicit SCMs with Soft Interventions
(Equivalence up to component-wise reparameterization) Let $\mathcal{M}=(\mathcal{A},\mathcal{X},g,\mathcal{I})$ and $\mathcal{M}^{\prime}=(\mathcal{A}^{\prime},\mathcal{X},g^{\prime},\mathcal{I})$ be two Latent Causal Models (LCM) based on AICMs $\mathcal{A},\mathcal{A}^{\prime}$ with shared observation space $\mathcal{X}$, shared intervention targets $\mathcal{I}$, and respective decoders $g$ and $g^{\prime}$. We say that $\mathcal{M}$ and $\mathcal{M}^{\prime}$ are equivalent up to component-wise reparameterization $\mathcal{M}\sim_{r}\mathcal{M}^{\prime}$ if there exists a component-wise transformation (Definition A1.2) $\phi_{\mathcal{Z}}$ from the causal variables $\mathcal{Z}$ to the causal variables $\mathcal{Z}^{\prime}$ and a component-wise transformation $\phi_{\mathcal{E}}$ between $\mathcal{E}$ and $\mathcal{E}^{\prime}$ such that:
1. Indices are preserved (i.e., $\phi_{i}(z_{i})=z^{\prime}_{i}$ and $\phi_{i}(e_{i})=e^{\prime}_{i}$). Corresponding edges are preserved (i.e., $\mathcal{Z}_{i}\rightarrow\mathcal{Z}_{j}$ holds in $\mathcal{G}$ iff $\mathcal{Z}^{\prime}_{i}\rightarrow\mathcal{Z}^{\prime}_{j}$ holds in $\mathcal{G}^{\prime}$. Edges $\mathcal{E}_{i}\rightarrow\mathcal{Z}_{i}$ should be preserved as well.)
2. The exogenous transformation preserves the probability measure on exogenous variables $p_{\mathcal{E}^{\prime}}=(\phi_{\mathcal{E}})_{*}p_{\mathcal{E}}$ (Definition A1.4).
3. The causal transformation preserves the probability measure on causal variables $p_{\mathcal{Z}^{\prime}}=(\phi_{\mathcal{Z}})_{*}p_{\mathcal{Z}}$ (Definition A1.4).
Theorem 3.5.
(Identifiability of latent causal models.) Let $\mathcal{M}=(\mathcal{A},\mathcal{X},g,\mathcal{I})$ and $\mathcal{M}^{\prime}=(\mathcal{A}^{\prime},\mathcal{X},g^{\prime},\mathcal{I})$ be two LCMs with shared observation space $\mathcal{X}$ and shared intervention targets $\mathcal{I}$. Suppose the following conditions are satisfied:
1. Data generating assumptions explained in Assumption 3.1.
2. Soft interventions satisfy Assumption 3.3.
3. The causal and exogenous variables are real-valued.
4. The causal and exogenous variables follow a multivariate normal distribution.
Then the following statements are equivalent:
-Two LCMs $\mathcal{M}$ and $\mathcal{M}^{\prime}$ assign the same likelihood to interventional and observational data i.e., $p^{\mathcal{X},\mathcal{I}}_{\mathcal{M}}(x,\tilde{x},i)=p^{\mathcal{X},\mathcal{I}}_{\mathcal{M}^{\prime}}(x,\tilde{x},i)$.
3.4 Training Objective
Consequently, there will be three latent variables in ICRL-SM:
1. A causal mechanism switch variable $\mathcal{V}$.
2. The pre-intervention exogenous variables $\mathcal{E}$.
3. The post-intervention exogenous variables $\tilde{\mathcal{E}}$.
@@ 3.4 Training Objective
1. $p(e)$ is the prior distribution over exogenous variables $e$.
2. $p(v)$ is the prior distribution over switch variables $v$.
3. $p(\tilde{e}|e,v)$ is a transition model that shows how the exogeneous variables change as a function of the intervention.
We factorize the posterior with a mean-field approximation $q(v,e,\tilde{e}|x,\tilde{x})=q(v|\tilde{x}-x)\cdot q(e|x)\cdot q(\tilde{e}|\tilde{x})$ and, following our data generation model (Figure 2(b)), the reconstruction probability as $p(x,\tilde{x}|e,\tilde{e},v)=p(x|e)p(\tilde{x}|\tilde{e})p(\tilde{x}-x|v)$. The prior over latent variables is factorized as $p(\tilde{e},e,v)=p(\tilde{e}|e,v)p(v)p(e)$(Figure 2(b)). Pre-intervention exogenous variables are mutually independent, hence, $p(e)=\Pi_{i}p(e_{i})$ and $p(v)=\Pi_{i}p(v_{i})$. We assume $p(e_{i})$ and $p(v_{i})$ to be standard Gaussian. Furthermore, as we assume $e_{i}=\tilde{e}_{i}$ for all non-intervened variables, the $p(\tilde{e}|e,v)$ will be as follows:
$p(\tilde{e}|e,v)=\Pi_{i\notin I}\delta(\tilde{e}_{i}-e_{i})\Pi_{i\in I}p(\tilde{e}_{i}|e,v)=\Pi_{i\notin I}\delta(\tilde{e}_{i}-e_{i})\Pi_{i\in I}p(\tilde{z}_{i}|e_{i})\left|\frac{\partial\tilde{z}_{i}}{\partial\tilde{e}_{i}}\right|$ (5)
The last equality is obtained from the Change of Variable Rule in probability theory, applied to the solution function $\tilde{z}_{i}=s_{i}(\tilde{e}_{i};e_{/i},h_{i}(v))$. Furthermore, we write $p(\tilde{z}_{i}|e,v)=p(\tilde{z}_{i}|e_{i})$ since only $e_{i}$ is a known parent of $\tilde{z}_{i}$ in implicit modeling. We assume $p(\tilde{z}_{i}|e_{i})$ to be a Gaussian whose mean is determined by $e_{i}$. We implement the solution function using a location-scale noise models [12] as also practiced in [3], which defines an invertible diffeomorphism. For simplicity, in our experiments, we are only going to change the $loc$ network in post-intervention. Therefore, $h_{i}(v)$ will be used as:
$\tilde{z}_{i}=\tilde{s_{i}}(\tilde{e}_{i};e_{/i},h_{i}(v))=\frac{\tilde{e}_{i}-(loc_{i}(e_{/i})+h_{i}(v))}{scale_{i}(e_{/i})},$ (6)
where $loc_{i}:\mathbb{R}^{n-1}\rightarrow\mathbb{R}$ and $scale_{i}:\mathbb{R}^{n-1}\rightarrow\mathbb{R}$ are fully connected networks calculating the first and second moments, respectively. The general overview of the model is illustrated in Figure 2(a).